OpenStudy (anonymous):

Help me with this weird question!! A tumor in a rat responds to a new radiation treatment by shrinking 5% with each treatment. How many treatments before the tumor has shrunk to a quarter of its original size? HINT: The size of the tumor doesn't matter.

OpenStudy (anonymous):

a) approximately 38 treatments b) approximately 11 treatments c) approximately 27 treatments d) approximately 75 treatments

OpenStudy (anonymous):

Well, a quarter is how much percent?

OpenStudy (anonymous):

25%

OpenStudy (anonymous):

So, if the treatment is increasing by 5% each time, it would 5 treatments, 5 x 5 is 25, right? But that's not one of the answers...

OpenStudy (anonymous):

oh no

OpenStudy (anonymous):

\[(0.5)^n = 0.25\] Find n.

OpenStudy (anonymous):

ack 0.05

OpenStudy (anonymous):

to decrease a number by \(5\%\) multiply it by \(100\%-5\%=95\%=.95\)

OpenStudy (anonymous):

it has to shrink 75% to get to a quarter..

OpenStudy (anonymous):

your job is to solve \[.95^x=.25\]

OpenStudy (anonymous):

Ack, yeah. Listen to satellite. Sorry. Misread the question slightly.

OpenStudy (anonymous):

I misread the question as well. \

OpenStudy (anonymous):

for this you need logarithms by the change of base formula if \[.95^x=.25\] then \[x=\frac{\ln(.25)}{\ln(.95)}\] and a calculator

OpenStudy (anonymous):

if you have not got to logs get, then you can guess and check which number \(x\) gives you \[.95^x=.25\]

OpenStudy (anonymous):

Ooooh!! 27?

OpenStudy (anonymous):

that is what i get, yes

OpenStudy (anonymous):

thank you all :)